OpenStudy (anonymous):

Which expression is equivalent to 1 – cos2 x?

OpenStudy (anonymous):

mmmnn Pythagorean identities? ^^

OpenStudy (anonymous):

do you know this..?? \[\sin ^2x+\cos ^2x=1\]

OpenStudy (anonymous):

then it will be simply as this \[\sin ^2x=1-\cos ^2x\]